# Investigating the Origin of CMB Large-Scale Features Using LiteBIRD and CMB-S4

### v1

@@ 4 Results of Data Analysis
Figure 2: [Top] Primordial power spectra of the best-fit candidates. The best-fit featureless candidate is represented by the black dashed line. [Bottom] CMB power spectrum residuals of the best-fit candidates with respect to the best-fit featureless candidate. The Planck measurements are represented by the gray points. For a better visualization, we bin $-\ell$ data ($\ell>30$) using a binwidth $\Delta\ell=60$.
@@ 4.1 Best Fit Candidates
Lastly, the best-fit candidate for the dark dimension model coincides with the featureless baseline model. [^5] One might wonder why the best fit value for $N_{\rm DD}-N_{*}$ is not $-6.6$, corresponding to the edge of our chosen prior. Since $N_{DD}-N_{*}=-6.43$ is nearly indistinguishable from a featureless model within observable scales (accounting for the cosmic variance at low $\ell$), decreasing $N_{DD}-N_{*}$ further only yields a negligible improvement in $\chi^{2}$. However, it is worth pointing out that, while the minimizer can get close to a global minimum, if the search is initiated from a different point in parameter space, there can be small differences in the converged values of the $\chi^{2}$. Therefore, this slight difference between $-6.6$ and $-6.43$ can be attributed to differences in the converged points by the minimizer. Nevertheless, for any small value of $N_{DD}-N_{*}$ which corresponds to a power law on cosmological scales, the best fit is $\Delta\chi^{2}\sim 0$ with respect to the baseline model. Any signature of this model within the observable scales provides a worse fit to the data since the angular power spectra from this model exhibit amplification at low multipoles.
@@ 5 Forecast Analysis
We assume one of our best-fit candidates to be the “fiducial,” i.e. the “true” model of the universe. We consider only the multiple range $\ell=2-2500$, corresponding to the scales probed by Planck. We simulate observations by adding noise to the fiducial according to the noise curves of a given combination of surveys. We use combinations of Planck data with LiteBIRD (PL + LB) and CMB-S4 (PL + LB + S4). We include CMB-S4 in our analysis to place additional constraints at small angular scales, while LiteBIRD will have the most constraining power at large scales. We assume PL to cover $\ell=800-2500$ with 20% sky coverage, LB to cover $\ell=2-39$ with 80% sky coverage and $\ell=40-1350$ with 20% sky coverage, and S4 to cover $\ell=40-2500$ with 40% sky coverage. These ranges ensure insignificant overlap between the surveys where their signal-to-noise ratios are comparable. The noise in the power spectrum for each survey is calculated as
@@
We fit each model to these simulated data using CosmoChord. We keep the same ranges on the priors for the cosmological and effective parameters as our Planck data analysis (see Table 1). We calculate the Bayes factor $\ln B=\ln\mathcal{Z}_{\rm fiducial}-\ln\mathcal{Z}_{\rm model}$ to assess the evidence for or against each model compared to the fiducial. A positive $\ln B$ means that the fiducial is favored. In the following discussion, we do not present results of forecasts for the dark dimension model since our data analysis has shown that such large-scale amplification is already severely constrained by the Planck data alone.
@@ 5.1 Results Assuming a Featureless Fiducial
Figure 3: Constraints on the step [top left], kink [top right], and $w-c_{s}^{2}$ [bottom left] models assuming a featureless fiducial. We mark the Planck best fit for each model with the small circles. Also shown are the projected Bayes factors [bottom right] obtained from our data and forecast analyses with respect to the featureless model.
@@
Figure 3 also shows the Bayes factors for these results. When adding the simulated measurements from LiteBIRD and CMB-S4, the kink model is strongly disfavored and the step model is substantially-to-strongly disfavored compared to the featureless fiducial. While the Bayes factors for the PL + LB + S4 forecasts between the step and kink models are consistent within $1\sigma$, the mean $\ln B$ for the step model is slightly smaller than that of the kink model. This could be because for the step model, two parameters can be adjusted to recover a featureless model — either the $\Delta N\rightarrow\infty$ limit or the $\Delta\mathcal{P}/\mathcal{P}_{0}\rightarrow 0$ limit can give a featureless power spectrum — while for the kink model, only the $\Delta\mathcal{P}/\mathcal{P}_{0}\rightarrow 0$ limit results in a featureless spectrum. For the $w-c_{s}^{2}$ model, there is a slight increase in the evidence against this model in the PL + LB forecast compared to the Planck analysis, but the evidence is still barely mentionable. When adding CMB-S4, the evidence against the $w-c_{s}^{2}$ model becomes substantial. Again, this is likely because CMB-S4 will be able to provide a better constraint on $w$ that is consistent with the $\Lambda$CDM model, making $c_{s}^{2}$ an irrelevant parameter and providing more evidence against this model. LSS measurements in conjunction with CMB-S4 would provide even better constraints on $w$.
@@ 5.2 Results Assuming a Fiducial with Features
Figure 5: Constraints on the the kink [top left] and $w-c_{s}^{2}$ [top right] models, assuming a step fiducial. We mark the Planck best fit of each model with the small circles. Also shown are the projected Bayes factors [bottom] obtained from our data and forecast analyses with respect to the step model.
@@
Figure 5 shows the constraints on the effective parameters for the kink and $w-c_{s}^{2}$ models from the Planck data analysis and the forecast analyses, assuming a step fiducial, as well as the Bayes factors for these results. Compared to our analysis with real Planck data, the parameters of the kink model are more constrained. In particular, despite $\Delta\mathcal{P}_{\rm kink}/\mathcal{P}_{0}$ being consistent with 00, we notice that its posterior peaks at a non-zero and negative value. The posterior of the parameter $N_{0}-N_{*}$ also displays a shift in the peak compared to the Planck data result, and smaller values of $\Delta N$ seem favored over larger ones. These are manifestations of the kink model trying to fit the large-scale fiducial dip. However, despite being slightly better than that for a featureless spectrum, the model is strongly disfavored over the step fiducial. Also, for the $w-c_{s}^{2}$ model, the results of the forecasts give similar results to the featureless fiducial — $w$ is more tightly constrained around $w=-1$ and $c_{s}^{2}$ is unconstrained for the PL + LB + S4 forecast. Again, this is most likely due to better constraints on the cosmological parameters, especially from CMB-S4.
@@
Primordial non-Gaussianities may help to distinguish between different large-scale primordial feature models (see [30] for a review), such as the step and kink models. To this end, we briefly discuss the primordial bispectra predicted by the best fit step and kink models. The bispectrum $B(k_{1},k_{2},k_{3})$ is defined as the three-point correlation of the curvature perturbation $\zeta$: $\langle\zeta_{\vec{k}_{1}}\zeta_{\vec{k}_{2}}\zeta_{\vec{k}_{3}}\rangle=(2\pi)^{3}\delta^{(3)}(\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3})B(k_{1},k_{2},k_{3})$. In Figure 6, we plot the dimensionless number $f_{\rm NL}(k)$ in the equilaterial limit $k_{1}=k_{2}=k_{3}=k$,
@@
for the step and kink best-fit candidates to the Planck data, where we calculated $B(k,k,k)$ and $\mathcal{P}_{0}(k)$ using BINGO. Note that, in this definition, we use the featureless power spectrum $\mathcal{P}_{0}(k)$ so that the features in $f_{\rm NL}(k)$ come directly from the bispectrum and are not introduced by the power spectrum. We can see that both the phases of the running and the amplitudes are different. However, it would be challenging to observationally constrain $f_{\rm NL}^{\rm feature}$ at such an order of magnitude.
@@ 6 Tests of $-\ell$ Anomalies Ignoring Short-Scale Effects
In our forecast analysis, we assumed that the true model of the universe was either featureless or contained a primordial step. In addition to the large dip that fits the $\ell\simeq 20$ anomaly in the $TT$ power spectrum, the step model also exhibits shorter-scale oscillations, and, importantly, the first of these oscillations following the dip has an amplitude which is (smaller but) comparable to the dip. This characteristic of the step model makes it testable across several multipoles, and our forecasts result in strong evidence in favor of the model (with the step fiducial) or against it (with the featureless fiducial). The kink model exhibits similar short-scale oscillations. It is, however, desirable to understand if future large-scale experiments, such as LiteBIRD, have the power to reveal the origin of the dip at $\ell\simeq 20$ and the suppression at $\ell\lesssim 10$ in the $TT$ data without appealing to detection of short-scale effects of the chosen model. To investigate this question, we consider two simplified diagnostic templates separately modeling the dip and suppression with no subsequent oscillations. While such templates are not motivated by realistic physics to our knowledge [^6] A burst of particle production can indeed produce a Gaussian bump feature in the primordial power spectrum [80, 81, 82], but not a dip. Such a bump is also followed by linear oscillations which are a signature of sharp features, but the envelope is such that their amplitude is very suppressed relative to the main bump. We could only find [83] in the literature in which a model displays a lognormal dip. However, it is unclear why that model does not exhibit the linear oscillations associated with sharp features., we consider these templates as a test of LiteBIRD’s capabilities to detect such features independent of short-scale effects.
@@ 6.1 Overview of the Diagnostic Templates
where $\Delta\mathcal{P}/\mathcal{P}_{0}$ is the maximum amplitude of the dip, $k_{0}$ is the dip center, and $\sigma$ is the feature width. Similarly, we model the $\ell\lesssim 10$ suppression using a primordal spectrum of the form
@@ 6.2 Results for the Dip Template
Figure 7: Results of the data and forecast analyses of the dip [left] and suppression [right] templates. The forecast analyses assume a dip and suppression fiducial, respectively. [Top] CMB power spectrum residuals of the best-fit candidates with respect to the best-fit featureless model. The Planck measurements are represented by the gray points. For a better visualization, we bin $-\ell$ data using a binwidth $\Delta\ell=60$. [Center] Constraints on the template parameters from our data and forecast analyses. [Bottom] Reconstructed primordial power spectra from the Planck data (gray colormap) and from our PL + LB + S4 forecasts (pink and teal colormaps). The best-fit featureless and template candidates are also plotted.
@@ 7 Conclusions
In this study, we investigated the prospects of observing $-\ell$ anomalies with future CMB temperature and anisotropy measurements. Since cosmic variance in the $TT$ power spectrum has already been reached by Planck, we focused on the ability of LiteBIRD to provide nearly cosmic-variance limited measurements of $E$ modes at large scales. These predicted measurements, in combination with the sharp $EE$ transfer function, can provide tight constraints on models that exhibit large-scale deviations from the standard cosmological model.
We analyzed four different classes of models that generate three types of large-scale signals in the $TT$ power spectrum: a primordial dip at $\ell\simeq 20$, primordial $\ell<10$ suppression, late-time $\ell<10$ suppression, and $\ell<10$ amplification. In the first part of our analysis, we summarized the agreement of our chosen models with the Planck temperature and $E$-mode data using Bayesian inference. We found that all models, except the large-scale amplification, can improve the fit to data by fitting the $-\ell$ anomalies, though none of them is statistically favored over the baseline model due to the extra number of parameters. Next, we performed a forecast of LiteBIRD and CMB-S4 measurements. We showed that, if the true model of the universe is the concordance model, we can strongly rule out a primordial large-scale dip and substantially-to-strongly rule out models predicting $\ell<10$ suppression. If the true model of the universe exhibits a dip at large scales, we predict strong evidence for LiteBIRD and CMB-S4 to detect such a feature. Moreover, these experiments will be able to strongly rule out primordial large-scale suppression and decisively rule out late-time large-scale suppression, assuming the large-scale dip as the true universe model. Lastly, we tested whether LiteBIRD can detect large-scale features while neglecting short-scale modifications to the CMB power spectrum, such as those imposed by sharp primordial features. Our results showed that LiteBIRD can decisively detect a large-scale dip but cannot detect large-scale suppression, without appealing to short-scale effects.

### v2

@@ 4 Results of data analysis
Figure 2: [Top] Primordial power spectra of the best-fit candidates. The best-fit featureless candidate is represented by the black dashed line. Note that the “best-fit" candidate for the DD model is an artifact from the boundary of priors (see the main texts and appendix for more details). [Bottom] CMB power spectrum residuals of the best-fit candidates with respect to the best-fit featureless candidate. The residuals of the Planck measurements with respect to the best-fit featureless candidate are represented by the gray points. For a better visualization, we bin $-\ell$ data ($\ell>30$) using a binwidth $\Delta\ell=60$.
@@ 4.1 Best fit candidates
Lastly, the “best-fit" candidate for the dark dimension model approaches to the featureless baseline model. This candidate is not at the location of the minimum $\chi^{2}$ which for this model is at the limit of the featureless baseline model, but is an artifact of the boundary of the chosen prior range. [^7] One might wonder why the best fit value for $N_{\rm DD}-N_{*}$ is not $-6.6$, corresponding to the edge of our chosen prior. Since $N_{DD}-N_{*}=-6.43$ is nearly indistinguishable from a featureless model within observable scales (accounting for the cosmic variance at low $\ell$), decreasing $N_{DD}-N_{*}$ further only yields a negligible improvement in $\chi^{2}$. However, it is worth pointing out that, while the minimizer can get close to a global minimum, if the search is initiated from a different point in parameter space, there can be small differences in the converged values of the $\chi^{2}$. The tolerance of $\chi^{2}$ is set to 0.1 in the search. Therefore, this slight difference of $\chi^{2}$ ($\sim-0.2$) between $N_{\rm DD}-N_{*}=-6.6$ and $N_{\rm DD}-N_{*}=-6.43$ can be attributed to differences in the converged points by the minimizer. Nevertheless, for any small value of $N_{DD}-N_{*}$ which corresponds to a power law on cosmological scales, the best fit is $\Delta\chi^{2}\sim 0$ with respect to the baseline model. We also tested that further lowering the lower limit of the $N_{DD}-N_{*}$ prior reduces its “best-fit” value as expected. This result can be readily understood: the signature of this model within the observable scales provides a worse fit to the data since the angular power spectra from this model exhibit amplification at low multipoles, but the current data prefers the opposite trend.
@@ 5 Forecast analysis
We assume one of our best-fit candidates to be the “fiducial,” i.e. the “true” model of the universe. We consider only the multiple range $\ell=2-2500$, corresponding to the scales probed by Planck. We simulate observations by adding noise to the fiducial according to the noise curves of a given combination of surveys. Note that these noise curves refer to the designs discussed in [44, 46], and the designs are expected to be constantly updated prior launch of the surveys. We use combinations of Planck data with LiteBIRD (PL + LB) and CMB-S4 (PL + LB + S4). We include CMB-S4 in our analysis to place additional constraints at small angular scales, while LiteBIRD will have the most constraining power at large scales. We assume PL to cover $\ell=800-2500$ with 20% sky coverage, LB to cover $\ell=2-39$ with 80% sky coverage and $\ell=40-1350$ with 20% sky coverage, and S4 to cover $\ell=40-2500$ with 40% sky coverage. These ranges ensure insignificant overlap between the surveys where their signal-to-noise ratios are comparable. The noise in the power spectrum for each survey is calculated as
@@
We fit each model to these simulated data using CosmoChord. We keep the same ranges on the priors for the cosmological and effective parameters as our Planck data analysis (see Table 1). We calculate the Bayes factor $\ln B=\ln\mathcal{Z}_{\rm fiducial}-\ln\mathcal{Z}_{\rm model}$ to assess the evidence for or against each model compared to the fiducial. A positive $\ln B$ means that the fiducial is favored, and we interpret the values of $\ln B$ using the Jeffreys’ scale [84]. In the following discussion, we do not present results of forecasts for the dark dimension model since our data analysis has shown that such large-scale amplification is already severely constrained by the Planck data alone.
@@ 5.1 Results assuming a featureless fiducial
Figure 3: Constraints on the step [top left], kink [top right], and $w-c_{s}^{2}$ [bottom] models assuming a featureless fiducial. We mark the Planck best fit for each model with the small circles.
Figure 4: The projected Bayes factors obtained from our data and forecast analyses with respect to the featureless model.
@@
Figure 4 shows the Bayes factors for these results. When adding the simulated measurements from LiteBIRD and CMB-S4, the kink model is strongly disfavored and the step model is substantially-to-strongly disfavored compared to the featureless fiducial. While the Bayes factors for the PL + LB + S4 forecasts between the step and kink models are consistent within $1\sigma$, the mean $\ln B$ for the step model is slightly smaller than that of the kink model. This could be because for the step model, two parameters can be adjusted to recover a featureless model — either the $\Delta N\rightarrow\infty$ limit or the $\Delta\mathcal{P}/\mathcal{P}_{0}\rightarrow 0$ limit can give a featureless power spectrum — while for the kink model, only the $\Delta\mathcal{P}/\mathcal{P}_{0}\rightarrow 0$ limit results in a featureless spectrum. For the $w-c_{s}^{2}$ model, there is a slight increase in the evidence against this model in the PL + LB forecast compared to the Planck analysis, but the evidence is still barely mentionable. When adding CMB-S4, the evidence against the $w-c_{s}^{2}$ model becomes substantial. Again, this is likely because CMB-S4 will be able to provide a better constraint on $w$ that is consistent with the $\Lambda$CDM model, making $c_{s}^{2}$ an irrelevant parameter and providing more evidence against this model. LSS measurements in conjunction with CMB-S4 would provide even better constraints on $w$.
@@ 5.2 Results assuming a fiducial with features
Figure 6: Constraints on the the kink [top left] and $w-c_{s}^{2}$ [top right] models, assuming a step fiducial. We mark the Planck best fit of each model with the small circles.
Figure 7: The projected Bayes factors obtained from our data and forecast analyses with respect to the step model.
@@
Figure 6 shows the constraints on the effective parameters for the kink and $w-c_{s}^{2}$ models from the Planck data analysis and the forecast analyses, assuming a step fiducial, and Figure 7 shows the Bayes factors for these results. Compared to our analysis with real Planck data, the parameters of the kink model are more constrained. In particular, despite $\Delta\mathcal{P}_{\rm kink}/\mathcal{P}_{0}$ being consistent with 00, we notice that its posterior peaks at a non-zero and negative value. The posterior of the parameter $N_{0}-N_{*}$ also displays a shift in the peak compared to the Planck data result, and smaller values of $\Delta N$ seem favored over larger ones. These are manifestations of the kink model trying to fit the large-scale fiducial dip. However, despite being slightly better than that for a featureless spectrum, the model is strongly disfavored over the step fiducial. Also, for the $w-c_{s}^{2}$ model, the results of the forecasts give similar results to the featureless fiducial — $w$ is more tightly constrained around $w=-1$ and $c_{s}^{2}$ is unconstrained for the PL + LB + S4 forecast. Again, this is most likely due to better constraints on the cosmological parameters, especially from CMB-S4.
@@
Primordial non-Gaussianities may help to distinguish between different large-scale primordial feature models (see [31] for a review), such as the step and kink models. To this end, we briefly discuss the primordial bispectra predicted by the best fit step and kink models. The bispectrum $B(k_{1},k_{2},k_{3})$ is defined as the three-point correlation of the curvature perturbation $\zeta$: $\langle\zeta_{\vec{k}_{1}}\zeta_{\vec{k}_{2}}\zeta_{\vec{k}_{3}}\rangle=(2\pi)^{3}\delta^{(3)}(\vec{k}_{1}+\vec{k}_{2}+\vec{k}_{3})B(k_{1},k_{2},k_{3})$. In Figure 8, we plot the dimensionless number $f_{\rm NL}(k)$ in the equilateral limit $k_{1}=k_{2}=k_{3}=k$,
@@
for the step and kink best-fit candidates to the Planck data, where we calculated $B(k,k,k)$ and $\mathcal{P}_{0}(k)$ using BINGO. Note that, in this definition, we use the featureless power spectrum $\mathcal{P}_{0}(k)$ so that the features in $f_{\rm NL}(k)$ come directly from the bispectrum and are not introduced by the power spectrum. We can see that both the phases of the running and the amplitudes are different. However, it would be challenging to observationally constrain $f_{\rm NL}^{\rm feature}$ at such an order of magnitude. We refer to [88, 89, 90, 91] for works and reviews on such data analyses.
@@ 6 Tests of $-\ell$ anomalies ignoring short-scale effects
In our forecast analysis, we assumed that the true model of the universe was either featureless or contained a primordial step. In addition to the large dip that fits the $\ell\simeq 20$ anomaly in the $TT$ power spectrum, the step model also exhibits shorter-scale oscillations up to $\ell\approx 250$, and, importantly, the first of these oscillations following the dip has an amplitude which is (smaller but) comparable to the dip. This characteristic of the step model makes it testable across several multipoles, and our forecasts result in strong evidence in favor of the model (with the step fiducial) or against it (with the featureless fiducial). The kink model exhibits similar short-scale oscillations. It is, however, desirable to understand if future large-scale experiments, such as LiteBIRD, have the power to reveal the origin of the dip at $\ell\simeq 20$ and the suppression at $\ell\lesssim 10$ in the $TT$ data without appealing to detection of short-scale effects of the chosen model. To investigate this question, we consider two simplified diagnostic templates separately modeling the dip and suppression with no subsequent oscillations. While such templates are not motivated by realistic physics to our knowledge [^8] A burst of particle production can indeed produce a Gaussian bump feature in the primordial power spectrum [92, 93, 94], but not a dip. Such a bump is also followed by linear oscillations which are a signature of sharp features, but the envelope is such that their amplitude is very suppressed relative to the main bump. We could only find [95] in the literature in which a model displays a lognormal dip. However, it is unclear why that model does not exhibit the linear oscillations associated with sharp features., we consider these templates in order to test LiteBIRD’s capabilities to detect such features independent of short-scale effects, rather than to analyze or constrain the templates themselves.
@@ 6.1 Overview of the diagnostic templates
where $\Delta\mathcal{P}/\mathcal{P}_{0}$ is the maximum amplitude of the dip, $k_{0}$ is the dip center, and $\sigma$ is the feature width. Similarly, we model the $\ell\lesssim 10$ suppression using a primordial spectrum of the form
@@ 6.2 Results for the dip template
Figure 9: CMB power spectrum residuals of the best-fit candidates for the dip and suppression templates with respect to the best-fit featureless model. The residuals of the Planck measurements with respect to the best-fit featureless candidate are represented by the gray points. For a better visualization, we bin $-\ell$ data using a binwidth $\Delta\ell=60$.
Figure 10: Results of the forecast analyses of the dip [left] and suppression [right] templates, assuming a dip and suppression fiducial, respectively. [Top] Constraints on the template parameters from our data and forecast analyses. [Bottom] Reconstructed primordial power spectra from the Planck data (gray colormap) and from our PL + LB + S4 forecasts (pink and teal colormaps). The best-fit featureless and template candidates are also plotted.
@@ 7 Conclusions
In this study, we investigated the prospects of observing $-\ell$ anomalies with future CMB temperature and polarization anisotropy measurements. Since cosmic variance in the $TT$ power spectrum has already been reached by Planck, we focused on the ability of LiteBIRD to provide nearly cosmic-variance limited measurements of $E$ modes at large scales. These predicted measurements, in combination with the sharp $EE$ transfer function, can provide tight constraints on models that exhibit large-scale deviations from the standard cosmological model.
We analyzed four different classes of models that generate three types of large-scale signals in the $TT$ power spectrum: a primordial dip at $\ell\simeq 20$, primordial $\ell<10$ suppression, late-time $\ell<10$ suppression, and $\ell<10$ amplification. In the first part of our analysis, we summarized the agreement of our chosen models with the Planck temperature and $E$-mode data using Bayesian inference. We found that all models, except the large-scale amplification, can improve the fit to data by fitting the $-\ell$ anomalies, though none of them is statistically favored over the baseline model due to the extra number of parameters. Next, we performed a forecast of LiteBIRD and CMB-S4 measurements. We showed that, if the true model of the universe is the concordance model, we can strongly rule out a primordial large-scale dip and substantially-to-strongly rule out models predicting $\ell<10$ suppression. If the true model of the universe exhibits a dip at large scales, we predict that LiteBIRD and CMB-S4 would find strong evidence in favor of detection. Moreover, these experiments would be able to strongly rule out primordial large-scale suppression and decisively rule out late-time large-scale suppression, assuming the large-scale dip as the true universe model. Lastly, we tested whether LiteBIRD can detect large-scale features while neglecting short-scale modifications to the CMB power spectrum, such as those imposed by sharp primordial features. Our results showed that LiteBIRD can decisively detect the large-scale dip but cannot detect the large-scale suppression, without appealing to short-scale effects.